I don't under stand how come i have to put 3x to the right and divide by 2 on this question, (3x + 2y = 6) and i really just don't understand how to do these
Found 2 solutions by stanbon, padmameesala:
Answer by stanbon(75887)

I don't under stand how come i have to put 3x to the right and divide by 2 on this question, (3x + 2y = 6) and i really just don't understand how to do these
------------
If you want to write the equation in slope-intercept
form you have to isolate "Y" on one side of the equal sign.
---------------------
Start: 3x + 2y = 6
Subtract 3x from both sides to get:
2y = -3x+6
Divide both sides by 2 to isolate "y":
y = (-3/2)x+3
==================
Cheers,
Stan H.

Answer by padmameesala(42)

3x + 2y =6
-3x -3x subtract 3x on both sides inorder to balance the equation
--------------
2y = 6 -3x
-- -----
2 2 divide by 2 on both sides inorder to balance the equation
y = (6-3x)/2
this is the way it is solved
hope you will understand
wish you good luck!